# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 5.2. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$
By applying the functor $\text{Ext}_{\mathcal{A}^{\vee}_{\mathbb{C}}}\left(B_{0}^{\mathbb{C}}(1)^{\otimes i}\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}-\right)$ to the cellular filtration of $B_{0}^{\mathbb{C}}(1)$, we may inductively compute these Ext groups by an aAHSS, where the base case was computed in Corollary 5.4. This spectral sequence takes the form
$E_{1}=\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}_{2}^{\mathbb{C}}\{1,\bar{\xi}_{1},\bar{\tau}_{1}\}\implies\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1)^{\otimes i+1}).$
The differentials take the same form as those in $\textbf{aAHSS}(B_{0}^{\mathbb{C}}(1))$, namely we have $d_{1}(\alpha[3])=h_{0}\alpha[2]$ and $d_{2}(\alpha[2])=h_{1}\alpha[0]$, where $\alpha\in\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$. For degree reasons, there is no $d_{3}$-differential. Hidden extensions may be obtained by comparison with the classical case, using the algebra map
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})\to\text{Ext}_{\mathcal{A}(1)^{\vee}}(B_{0}(1)^{\otimes i}).$
@@
5.3. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ and the ring of cooperations
Next, we compute $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ for all $k\geq 1$.
Theorem 5.6.
[CQ21, Theorem 3.38] There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{C}}$-comodules and $\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules:
$\frac{\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))}{v_{1}\textup{-torsion}}\cong\Sigma^{4k-4,2k-2}Z^{\mathbb{C}}_{\alpha(k)}\oplus\bigoplus_{j=0}^{4k-8}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}],$
@@ 5.3. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ and the ring of cooperations
(5.7) $0\to\Sigma^{2k,k}B_{0}^{\mathbb{C}}(\tfrac{k}{2})\to B_{0}^{\mathbb{C}}(k)\to B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\to 0$
Applying the functor $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(-)$ gives a long exact sequence of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules. Moreover, after killing $v_{1}$-torsion the connecting homomorphism is trivial, giving short exact sequences whose middle term is $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$. Therefore, this Ext group decomposes into the Ext groups of the kernel and cokernel of the original short exact sequence (5.7). The kernel is handled by the inductive hypothesis: we have
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\Sigma^{2k,k}B_{0}^{\mathbb{C}}(\tfrac{k}{2}))\cong\Sigma^{2k,k}\left(\Sigma^{2k-4,k-2}Z_{\mathbb{\alpha}(k/2)}^{\mathbb{C}}\oplus\bigoplus_{j=0}^{2k-8}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\right).$
@@
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\right)\cong\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1)).$
@@
$\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1))\cong\bigoplus_{j=0}^{k/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}].$
@@
(5.8) $0\to\Sigma^{2(k-1),k-1}B_{0}^{\mathbb{C}}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}B_{0}^{\mathbb{C}}(1)\to B_{0}^{\mathbb{C}}(k)\to B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\to 0$
Applying the functor $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(-)$ gives a long exact sequence of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules, and as in the even case, the connecting homomorphism is trivial modulo $v_{1}$-torsion. Thus, the Ext group we are interested in decomposes into the Ext groups of the kernel and cokernel of 5.8. In this case, the cokernel is simpler to compute. A change of rings isomorphism gives
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\right)\cong\text{Ext}_{\mathcal{A}(0)_{\mathbb{C}}^{\vee}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1)).$
@@
$\text{Ext}_{\mathcal{A}(0)_{\mathbb{C}}^{\vee}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1))\cong\bigoplus_{j=0}^{(k-1)/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}].$
@@
$E_{1}=\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2}))\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}_{2}^{\mathbb{C}}\{1,\bar{\xi}_{1},\bar{\tau}_{1}\}\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}B_{0}^{\mathbb{C}}(1)\right).$
By induction, we may rewrite $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2}))$ as
$\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2}))\cong\Sigma^{2k-6,k-3}Z_{\alpha(k)-1}^{\mathbb{C}}\oplus\bigoplus_{j=0}^{2k-10}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}],$
using that $\alpha(\tfrac{k-1}{2})=\alpha(k)-1.$ This splitting leads to a splitting of the $E_{1}$-page of the algebraic Atiyah–Hirzebruch spectral sequence, and we may analyze each summand individually. On the left-hand summand, Proposition 5.5 implies that result of the spectral sequence is isomorphic to $\Sigma^{2k-2,k-1}Z^{\mathbb{C}}_{\alpha(k)}$. On the right-hand summand, the spectral sequence collapses on the $E_{1}$-page, giving back the original summand modulo $v_{1}$-torsion. Thus, after reindexing the $h_{0}$-towers, the Ext groups of the kernel and cokernel and right-hand side of the Ext groups assemble to give the result.
@@
The above proof is where our argument differs most from the one given in [CQ21]. We explain these differences here. Classically, Mahowald [Mah81] showed that the groups $\text{Ext}_{\mathcal{A}(1)^{\vee}}(B_{0}(k))$ can be expressed solely in terms of bo and bsp. To be precise, for a spectrum $X$, let $X^{\langle n\rangle}$ denote the $n$-th Adams cover of $X$, so that
$X\leftarrow X^{\langle 1\rangle}\leftarrow X^{\langle 2\rangle}\leftarrow X^{\langle 3\rangle}\leftarrow\cdots$
is a minimal Adams resolution of $X$. Then there are isomorphisms:
$\frac{\text{Ext}_{\mathcal{A}(1)^{\vee}}(B_{0}(k))}{v_{1}\text{-torsion}}\cong\left\{\begin{array}[]{ll}\text{Ext}_{\mathcal{A}(1)^{\vee}}(\text{H}_{**}(\text{bo}^{\langle 2k-\alpha(k)\rangle}))&k\equiv 0\,(2),\\ \text{Ext}_{\mathcal{A}(1)^{\vee}}(\text{H}_{**}(\text{bsp}^{\langle 2k-\alpha(k)-1\rangle}))&k\equiv 1\,(2).\end{array}\right.$
One thing to note is that in the classical case, each of the above groups is describable by a direct sum of $h_{0}$-towers followed by some shift of $\text{Ext}_{\mathcal{A}(1)^{\vee}}(\mathbb{F}_{2})$. In particular, if $x\in\text{Ext}^{s,f}_{\mathcal{A}(1)^{\vee}}(B_{0}(k))$ is in stem $s<4(2k-\alpha(k))$, then $x$ does not support $h_{1}$-multiplication.
@@
$\frac{\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))}{v_{1}\text{-torsion}}\cong Z^{\mathbb{C}}_{2k-\alpha(k)}.$
However, this is not quite correct. The group $Z_{2k-\alpha(k)}^{\mathbb{C}}$ has classes in small stems which support $h_{1}$-multiplication, while the decomposition given by the short exact sequences of Proposition 2.21 do not allow for this. For example, for $k=2$ we have $32k-\alpha(k)=4-1=$ 3, hence Culver–Quigley claim that $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(2))$ is isomorphic to
$Z_{3}^{\mathbb{C}}=\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{4,2}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{5,3}\mathbb{M}_{2}^{\mathbb{C}}[h_{1}]/(h_{1}^{2})\oplus\Sigma^{8,4}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))\langle 2\rangle.$
We see that there is a class in stem 5 which supports an $h_{1}$-multiplication. Alternatively, using the short exact sequence of Proposition 2.21 we see that the group $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(2))$ is isomorphic to following the direct sum:
$\Sigma^{8,4}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))\oplus\text{Ext}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(0))\cong\Sigma^{8,4}Z_{1}^{\mathbb{C}}\oplus\mathbb{M}_{2}^{\mathbb{C}}[h_{0}],$
which agrees with the description in Proposition 5.5. This decomposition does not admit any $h_{1}$-multiplication in stem 5; in fact, there are no classes in stem 5. This type of difference extends to all values of $k\geq 2$, showing that the formula given in [CQ21, Theorem3.38] is incorrect. However, this observation is actually in line with the rest of the work done in Culver–Quigley and doesn’t affect any end results. More precisely:

### v2

@@ 5.2. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$
By applying the functor $\text{Ext}^{***}_{\mathcal{A}^{\vee}_{\mathbb{C}}}\left(B_{0}^{\mathbb{C}}(1)^{\otimes i}\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}-\right)$ to (4.3), we may inductively compute these Ext groups by an aAHSS, where the base case was computed in Corollary 5.3. This spectral sequence takes the form
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}_{2}^{\mathbb{C}}\{[1],]\bar{\xi}_{1}],[\bar{\tau}_{1}]\}\implies\text{Ext}_{\mathcal{A}(1)_{\mathbb{C}}^{\vee}}(B_{0}^{\mathbb{C}}(1)^{\otimes i+1}).$
The differentials take the same form as those in $\textbf{aAHSS}(B_{0}^{\mathbb{C}}(1))$, namely we have $d_{1}(\alpha[3])=h_{0}\alpha[2]$ and $d_{2}(\alpha[2])=h_{1}\alpha[0]$, where $\alpha\in\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})$. For degree reasons, there is no $d_{3}$-differential. Hidden extensions may be obtained by comparison with the classical case, using the algebra map
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1)^{\otimes i})\to\text{Ext}^{**}_{\mathcal{A}(1)^{\vee}}(B_{0}(1)^{\otimes i}).$
@@
5.3. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ and the ring of cooperations
Next, we compute $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$ for all $k\geq 1$.
Theorem 5.5 (Corrected from [17, Theorem 3.38]).
There is an isomorphism of $\mathcal{A}(1)^{\vee}_{\mathbb{C}}$-comodules and $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules:
$\frac{\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))}{v_{1}\textup{-torsion}}\cong\Sigma^{4k-4,2k-2}Z^{\mathbb{C}}_{\alpha(k)}\oplus\bigoplus_{j=0}^{k-2}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}],$
@@
(5.2) $0\to\Sigma^{2k,k}B_{0}^{\mathbb{C}}(\tfrac{k}{2})\to B_{0}^{\mathbb{C}}(k)\to B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\to 0.$
Applying the functor $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(-)$ gives a long exact sequence of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules. Moreover, after killing $v_{1}$-torsion the connecting homomorphism is trivial, giving short exact sequences whose middle term is $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))$. Therefore, this Ext group decomposes into the Ext groups of the kernel and cokernel of the original short exact sequence (5.2). The kernel is handled by the inductive hypothesis: we have
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\Sigma^{2k,k}B_{0}^{\mathbb{C}}(\tfrac{k}{2}))\cong\Sigma^{2k,k}\left(\Sigma^{2k-4,k-2}Z_{\mathbb{\alpha}(k/2)}^{\mathbb{C}}\oplus\bigoplus_{j=0}^{2k-8}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\right).$
@@
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\right)\cong\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1)).$
@@
$\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(\tfrac{k}{2}-1))\cong\bigoplus_{j=0}^{k/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}].$
@@
(5.3) $0\to\Sigma^{2(k-1),k-1}B_{0}^{\mathbb{C}}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}B_{0}^{\mathbb{C}}(1)\to B_{0}^{\mathbb{C}}(k)\to B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\to 0.$
Applying the functor $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(-)$ gives a long exact sequence of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(\mathbb{M}_{2}^{\mathbb{C}})$-modules, and as in the even case, the connecting homomorphism is trivial modulo $v_{1}$-torsion. Thus, the Ext group we are interested in decomposes into the Ext groups of the kernel and cokernel of (5.3). In this case, the cokernel is simpler to compute. A change of rings isomorphism gives
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1)\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))_{\mathbb{C}}^{\vee}\right)\cong\text{Ext}^{***}_{\mathcal{A}(0)_{\mathbb{C}}^{\vee}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1)).$
@@
$\text{Ext}^{***}_{\mathcal{A}(0)_{\mathbb{C}}^{\vee}}(B_{1}^{\mathbb{C}}(\tfrac{k-1}{2}-1))\cong\bigoplus_{j=0}^{(k-1)/2-1}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}].$
@@
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2}))\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}\mathbb{M}_{2}^{\mathbb{C}}\{[1],[\bar{\xi}_{1}],[\bar{\tau}_{1}]\}\implies\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}\left(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2})\otimes_{\mathbb{M}_{2}^{\mathbb{C}}}B_{0}^{\mathbb{C}}(1)\right).$
By induction, we may rewrite $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2}))$ as
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(\tfrac{k-1}{2}))\cong\Sigma^{2k-6,k-3}Z_{\alpha(k)-1}^{\mathbb{C}}\oplus\bigoplus_{j=0}^{2k-10}\Sigma^{4j,2j}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}],$
using that $\alpha(\tfrac{k-1}{2})=\alpha(k)-1.$ This splitting leads to a splitting of the $\textup{E}_{1}$-page of the algebraic Atiyah–Hirzebruch spectral sequence, and we may analyze each summand individually. On the left-hand summand, Proposition 5.4 implies that result of the spectral sequence is isomorphic to $\Sigma^{2k-2,k-1}Z^{\mathbb{C}}_{\alpha(k)}$. On the right-hand summand, the spectral sequence collapses on the $\textup{E}_{1}$-page, giving back the original summand modulo $v_{1}$-torsion. Thus, after reindexing the $h_{0}$-towers, the Ext groups of the kernel and cokernel and right-hand side of the Ext groups assemble to give the result.
@@
The above proof is where our argument differs most from the one given in [17]. We explain these differences here.
@@
$\frac{\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(k))}{v_{1}\text{-torsion}}\cong Z^{\mathbb{C}}_{2k-\alpha(k)}.$
However, this is not quite correct. Let us illustrate this difference in the case of $B_{0}^{\mathbb{C}}(2)$.
By Proposition 2.11, there is a short exact sequence of $\mathcal{A}(1)^{\vee}_{\mathbb{C}}$-comodules
(5.5) $0\to\Sigma^{4,2}B_{0}^{\mathbb{C}}(1)\to B_{0}^{\mathbb{C}}(2)\to B_{1}^{\mathbb{C}}(0)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))^{\vee}_{\mathbb{C}}\to 0.$
Applying the functor $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}^{***}(-)$ gives a long exact sequence of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}^{***}(\mathbb{M}_{2}^{\mathbb{C}})$-modules. The Ext group of the kernel of (5.5) has already been explicitly determined (see Figure 5.3). The Ext group of the cokernel may be written using a change of rings isomorphism as
$\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(0)\otimes(\mathcal{A}(1)\operatorname{/\!/}\mathcal{A}(0))^{\vee}_{\mathbb{C}})\cong\text{Ext}^{***}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(0))\cong\mathbb{M}_{2}^{\mathbb{C}}[h_{0}].$
For degree reasons, the connecting homomorphism
$\text{Ext}^{s,f,w}_{\mathcal{A}(0)^{\vee}_{\mathbb{C}}}(B_{1}^{\mathbb{C}}(0))\xrightarrow{\delta}\text{Ext}^{s-1,f,w}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))$
is trivial, implying that the long exact sequence in Ext decomposes into short exact sequences. Thus, we have an isomorphism
$\text{Ext}_{\mathcal{A}(1)^{\vee}}^{***}(B_{0}^{\mathbb{C}}(2))\cong\Sigma^{4,2}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))\oplus\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\cong\Sigma^{4,2}Z_{1}^{\mathbb{C}}\oplus\mathbb{M}_{2}^{\mathbb{C}}[h_{0}],$
which agrees with 5.5 as $\alpha(2)=1$. We depict this group in Figure 5.4.
Figure 5.4. The group $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}^{***}(B_{0}^{\mathbb{C}}(2))$.
However, Culver–Quigley claim that $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(2))$ is isomorphic to
$Z_{3}^{\mathbb{C}}=\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{4,2}\mathbb{M}_{2}^{\mathbb{C}}[h_{0}]\oplus\Sigma^{5,3}\mathbb{M}_{2}^{\mathbb{C}}[h_{1}]/(h_{1}^{2})\oplus\Sigma^{8,4}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{C}}}(B_{0}^{\mathbb{C}}(1))\langle 2\rangle.$
There is a class in stem 5 which supports an $h_{1}$-multiplication. We depict this group in Figure 5.5.
Figure 5.5. The group $Z_{3}^{\mathbb{C}}$.
This type of difference extends to all values of $k\geq 2$, showing that the formula given in [17, Theorem 3.38] is incorrect. However, this observation is actually in line with the rest of the work done in Culver–Quigley and doesn’t affect any end results. More precisely: